OpenStudy (anonymous):

A little confused on how to do this one... Solve the system by substitution. { 2x - y + z = -4 { z = 5 { -2x + 3y - z = -10 Help please. :P

OpenStudy (texaschic101):

2x - y + z = -4 .....now sub in 5 for z 2x - y + 5 = -4 2x - y = -4 - 5 2x - y = -9 2x + 3y - z = -10 ....sub in 5 for z 2x + 3y - 5 = -10 2x + 3y = -10 + 5 2x + 3y = -5 now we have : 2x - y = -9 --- multiply by -1 2x + 3y = -5 ------------ -2x + y = 9 (result of multiplying by -1) 2x + 3y = -5 ------------add 4y = 4 y = 1 now use an original equation and sub in known values to find x 2x - y + z = -4 2x - 1 + 5 = -4 2x + 4 = -4 2x = -4 - 4 2x = -8 x = -4 check.. 2x + 3y - z = -10 2(-4) + 3(1) - 5 = -10 -8 + 3 - 5 = -10 -13 + 3 = -10 -10 = -10 (correct) x = -4, y = 1, and z = 5